60 ml of mixture of CO and H2 were mixed with 40 ml of oxygen. Mixture was exploded and after cooling volume was found to be 30 ml. Find the volume of CO and H2. (If all the volume are measured at STP).

Grade:11

1 Answers

varun
6 years ago
Hello kumar brahma
Let the volume of  mixture of CO be a ml
the volume of mixture be b ml
therefore a+b=60ml
Now CO + ½ H2 gives CO2
and then H2 + ½ O2 gives H2O
Afer cooling,the volume is found to be 30ml
therefore
a+{40-(a/2 + b/2)} = 30
afer solving you get
10=b/2 – a/2
b-a=20 ml
therefore
b=40 ml and a=20 ml
Therefore volume of CO=20 ml
and volume of H2=40ml
THANK YOU